a company buys the following 11 computers at $1350 each, 2 printers at $2600 each 60 Boxes of paper at $150 per dozen boxes. what is the total cost of this purchase? can u please explain how u got the answer please thank u.
Answer by user_dude2008(1862)

11*1350+2*2600+60*150 = 14850+5200+9000 = 29050

total cost is $29,050